the ratio of the number of boys in a class to the number of girls is 11 to 10. if there are 525 students in the class, how many are boys? Girls?
Answer by richwmiller(17219)

11x are boys
10x are girls
21x = total students
21x=525
x=525/21
x=25
10*25=250 girls
11*25=275 boys